Hannah received an estate of $50,000 from her grandmother. She invested it and received 12 percent interest compounded annually.
(a) Write an equation that expresses the amount of money in the account as a function of n, where n equals the numbers of years since the original deposit was made.
(b) Use the equation in (a) to determine how much money is in the account at the end of n = 10 years.
(c) What was the total interest earned at n = 10 years?
Write the equation for (a) on your written work and answer parts (b) and (c).

Answer by mananth(16946)


A = P*(1+r)^nt
A = amount aftern years
P= Principal
n= number of years
t = number of compounding periods
r= rate of iterest
I = interest
A =50000*(1+0.12)^10*1
A= 50000*(1.12)^10
A=$155.292.41
A-P = I
155.292.41-50000= $105,292.41